OpenStudy (anonymous):

C=5p + 215 solve for P Would it be (c-215)/5 = P or c/5 - 215 = P

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

the first one :-) but if you wanted you could rewrite it: (c-215)/5=P c/ 5 - 215/5 = p c/5 - 43 =P Usually though you are expected to have a common denominator :-)

OpenStudy (anonymous):

how do you know when to divide by just the first variable or both?

OpenStudy (anonymous):

Always divide the whole side of the equation.

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

You can later distribute things our properly.